OpenStudy (anonymous):

A video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. On level 1, a player must earn 1 experience point. On level 2, a player must earn 7 experience points. On level 3, a player must earn 15 experience points. On level 4, a player must earn 25 experience points. Which recursive equation represents the pattern? Multiple Choice: an = an – 1 + n an = an – 1 + 3n + 1 an = an – 1 + 5n an = an – 1 + 2(n + 1)

OpenStudy (anonymous):

i don't know lets check on how about \( a_n = a_{n – 1} + 2(n + 1)\) and maybe we will get lucky

OpenStudy (anonymous):

im not quite sure what to do

OpenStudy (anonymous):

\(a_1=1\) \(a_2=7=^?1+2(2+1)=^?1+6=7\) ok that one works for \(a_2\) lets see if it continues to work

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\(a_3=15=^?7+2(3+1)=^?7+8=15\) ok that one looks good too!

OpenStudy (anonymous):

\(a_4=25=15+2\times 5=15+10=25\) whew i would have been annoyed if we had to try a different formula looks like the last one works, just by trying it

OpenStudy (anonymous):

Okay thanks can you help me with another one please